# &Jiawei Li

### v1

@@ Abstract
Addressing data integrity challenges, such as unlearning the effects of data poisoning after model training, is necessary for the reliable deployment of machine learning models. State-of-the-art influence functions, such as EK-FAC (Grosse et al., 2023), often fail to accurately attribute abnormal model behavior to the specific poisoned training data responsible for the data poisoning attack. In addition, traditional unlearning algorithms often struggle to effectively remove the influence of poisoned samples (Pawelczyk et al., 2024), particularly when only a few affected examples can be identified (Goel et al., 2024). To address these challenge, we introduce $\Delta$-Influence, a novel approach that leverages influence functions to trace abnormal model behavior back to the responsible poisoned training data using as little as just one poisoned test example. $\Delta$-Influence applies data transformations that sever the link between poisoned training data and compromised test points without significantly affecting clean data. This allows $\Delta$-Influence to detect large negative shifts in influence scores following data transformations, a phenomenon we term as influence collapse, thereby accurately identifying poisoned training data. Unlearning this subset, e.g. through retraining, effectively eliminates the data poisoning. We validate our method across three vision-based poisoning attacks and three datasets, benchmarking against four detection algorithms and five unlearning strategies. We show that $\Delta$-Influence consistently achieves the best unlearning across all settings, showing the promise of influence functions for corrective unlearning. Our code is publicly available at: https://github.com/andyisokay/delta-influence
@@ 1 Introduction
Machine learning models are increasingly deployed in critical sectors such as healthcare, finance, and autonomous systems (Davenport and Kalakota, 2019; Huang et al., 2020; Chatila et al., 2021; Soori et al., 2023). This widespread adoption underscores the importance of ensuring model integrity and robustness against malicious attacks, particularly data poisoning attacks. In data poisoning, adversaries intentionally manipulate training data by introducing carefully crafted, often imperceptible modifications (Chatila et al., 2021), leading to incorrect predictions or embedding specific malicious behaviors within the trained models (Fan et al., 2022). Given the large scale of modern datasets, identifying and removing all manipulated samples is typically impractical (Nguyen et al., 2024a; Goel et al., 2024). Therefore, a viable approach involves detecting and attributing the impact of data poisoning to a small set of influential training data points, which is unlearned to mitigate the data poisoning attack.
@@
The challenge of effective unlearning largely depends on the extent of knowledge about the data poisoning attack. For example, Goel et al. (2024) demonstrate that retraining a model after removing a randomly sampled subset containing half of the manipulated data fails to eliminate poisoning in relatively simple attacks like BadNet (Gu et al., 2019). In contrast, retraining without the entire set of manipulated data successfully removes the attack, highlighting the limitations of partially identifying manipulated data. Furthermore, for more sophisticated poisoning strategies such as Witches’ Brew (Geiping et al., 2021), Pawelczyk et al. (2024) reveal that existing unlearning algorithms are ineffective unless the model is retrained without the full manipulated set, even when full access to the manipulated data is available.
Building upon the framework of Corrective Unlearning introduced by Goel et al. (2024), our work addresses a real-world scenario in which the defender has identified a small set of affected test points. In this work, we address the hardest of such settings where only a single test point affected by the data poisoning attack is identifed. Detecting such an affected test point is a natural prerequisite for initiating the unlearning process. This detection either occurs naturally during deployment due to the occurrence of abnormal predictions or via exhaustive in-house stress testing. Leveraging this poisoned test point, our approach comprises two primary tasks: first, identifying a critical set of manipulated training points responsible for the compromised prediction; and second, applying unlearning algorithms (Goel et al., 2023; Kurmanji et al., 2023; Foster et al., 2024) to remove the influence of these points from the model.
@@
Within this framework, influence functions (Koh and Liang, 2017b) serve as a natural tool for attributing model predictions to specific training data points. However, recent studies (Grosse et al., 2023; Nguyen et al., 2024b; Bae et al., 2024) have indicated that state-of-the-art influence functions struggle to accurately identify the manipulated data points in deep learning models when used in a naive manner. Our experiments in Section 3 also corroborate this finding. To address this limitation, we introduce $\Delta-\mathrm{Influence}$, a novel approach that enhances influence functions to reliably identify a sufficient critical set of training data points necessary for unlearning data poisoning without compromising model accuracy. Instead of directly calculating each training point’s influence on a poisoned test point, $\Delta-\mathrm{Influence}$ assesses the change in influence scores before and after perturbing the test point through label flipping and image transformation. As shown using ablation studies in Section 4, label flipping is essential for breaking the association between poisoned data and the compromised test point, while image transformations introduces randomness that reduces false positive rates by preserving the influence of benign data. These combined perturbations improve the detection of poisoned training samples.
@@
To assess the effectiveness of $\Delta-\mathrm{Influence}$ and the broader applicability of influence functions in this context, we apply our method to three prominent data poisoning attacks: BadNet (Gu et al., 2019), Witches’ Brew (Geiping et al., 2021), and Smooth Trigger (Zeng et al., 2021). We compare our approach against multiple defenses (Chen et al., 2018; Tran et al., 2018; Zeng et al., 2021; Grosse et al., 2023) that operate with similar or less information about the poisoning than $\Delta-\mathrm{Influence}$. Each attack presents unique challenges for detection and mitigation, as evidenced by the varying performance of existing detection methods across different attacks. Additionally, we conduct experiments using several known unlearning algorithms to remove the data poisoning using the identified set. These experiments provides a comprehensive comparison of these unlearning algorithms highlight the strengths and weaknesses of each. For example, gradient ascent-based methods like SCRUB (Kurmanji et al., 2023) and weight deletion methods like SSD (Foster et al., 2024) can effectively unlearn poisoning when the detected set of training poisons is reasonably accurate. However, their resultant accuracy drops significantly if the detected set includes many falsely flagged clean examples. In contrast, methods like EU and CF (Goel et al., 2024) are surprisingly robust to false positives, delivering the best unlearning and accuracy. Overall, our experiments demonstrate that $\Delta-\mathrm{Influence}$ consistently outperforms existing algorithms across all settings, offering a robust defense against sophisticated data poisoning attacks while preserving accuracy.
@@ 2 Using Influence functions to detect poisons
However, our experiments in Section 3, along with several recent studies (Nguyen et al., 2024b; Bae et al., 2024; Li et al., 2024b), demonstrate that naively applying state-of-the-art influence functions (Grosse et al., 2023) fails to accurately identify poisoned points in deep neural networks. This limitation necessitates the development of a more robust method to effectively utilize influence functions for detecting and unlearning data poisoning.
@@ 2.1 Our Algorithm: $\Delta$-Influence
To address the shortcomings of the above naive approach, we introduce $\Delta-\mathrm{Influence}$, a novel algorithm for influence-function based detection of poisoned training samples. The core idea of $\Delta-\mathrm{Influence}$ is to monitor the changes in influence scores of training data points when the affected test point undergoes various transformations.
@@
Formally, let $g_{j}$ be a transformation applied to the test point $z_{\text{te}}=\left(x_{\text{te}},y_{\text{te}}\right)$, consisting of pairing the test image with a random label $y_{\text{te}}^{\prime}$ and applying standard data augmentations such as blurring, color jitter, rotating, and flipping to $x_{\text{te}}$ (see Appendix B for a list of all transformations). In this work, we utilize common data augmentation techniques without designing any poison-specific transformations. We consider this simplicity of our approach and not needing to design special transformations to be a key strength of our work. Then, for each transformation $g_{j}$, compute the change in influence score as:
@@
This is illustrated in Figure 2, where $\Delta\mathrm{Infl}({i},{j})$ is consistently negative for poisoned samples across all transformations, whereas it often remains near zero (compared to that of poisons) or shows no clear trend for clean examples. However, Figure 2 shows that this is not consistently the case for all clean examples (with some values being considerably small), which brings us to the next component.
Boosting Using Multiple Transformations. The above discussion shows that the change in influence score $\Delta\mathrm{Infl}({i},{j})$ can be used as a score function for detecting whether $z_{\text{tr}}^{i}$ is a manipulated training sample. However, this score function is a relatively weak classifier, especially for clean points, as seen in Figure 2. To overcome this problem, we use classical ideas from bagging and apply multiple transformations $g_{1},\ldots,g_{n_{b}}$ to obtain a series of weak score functions. Specifically, we use $n_{b}$ transformations to obtain $n_{b}$ weak score functions.
@@
Unlearning identified points. Once the set of poisoned training points $\mathcal{P}$ is identified using $\Delta-\mathrm{Influence}$, the next step is to unlearn these points to mitigate the data poisoning attack. We employ several unlearning algorithms (Goel et al., 2023; Kurmanji et al., 2023; Golatkar et al., 2020; Foster et al., 2024) to remove the influence of $\mathcal{P}$ from the trained model $\theta^{\star}$. In practice, the choice of unlearning algorithm may depend on factors such as computational efficiency, scalability, and the specific characteristics of the poisoning attack. In this work, we look at several popular algorithms including retraining from scratch (denoted as EU (Goel et al., 2023)), CF (Goel et al., 2023), SSD (Foster et al., 2024), SCRUB (Golatkar et al., 2020), and BadT (Kurmanji et al., 2023).
@@ 2.2 Full Algorithm
    Influence Score For each training data point $z_{\text{tr}}^{i}\in\mathcal{D}$ and each transformation $g_{j}\in\mathcal{G}$, compute the change in influence score $\Delta\mathrm{Infl}({i},{j})$ as defined in Equation 2.
@@ 3.1 Experimental Setup
    Patch Trigger (BadNet)  (Gu et al., 2019): Also studied in Goel et al. (2024), this attack involves adding a small, subtle patch to the corner of selected training images and altering their labels to a designated target class. The presence of the patch causes the trained model to misclassify any test image containing the patch into the target class while maintaining normal performance on other inputs.
@@
    Frequency Trigger  (Zeng et al., 2021): In this approach, a trained, imperceptible pattern is embedded both the spatial and frequency domains, thereby encompassing the whole image unlike BadNet. As shown in Alex et al. (2024). these patterns are difficult to detect by both human observers and automated detection methods, making the poisoned samples challenging to identify and remove.
@@
    Clean Label Attack (Witches’ Brew)  (Geiping et al., 2021): Unlike BadNets and Frequency Trigger, this attack adds an imperceptible pattern to images without altering their labels. The poisoned samples appear benign since their labels are consistent with their content, yet they cause the model to learn incorrect associations, leading to misclassifications during inference. As shown in Pawelczyk et al. (2024), these patterns are difficult to unlearn using unlearning algorithms.
Model and Datasets. We utilize the CIFAR10 and CIFAR100 datasets  (Krizhevsky, ) and a ResNet18 model (He et al., 2015), following the standard benchmarks and models used in the state-of-the-art machine unlearning setup (Pawelczyk et al., 2024). For CIFAR10, we poison 500 training images (1% of the dataset), while for CIFAR100, we poison 125 training images (0.25% of the dataset) for all attack types except BadNet, which requires a higher size of 350 samples to be effective. The victim class and attack class (when different) are selected randomly. Detection methods are tuned on a small validation set using cross-validation techniques. Hyperparameters such as threshold values and clustering parameters are optimized based on validation performance metrics to achieve the best balance between detection accuracy and false positive rates. Detailed hyperparameter settings are provided in Appendix B to ensure reproducibility.
Compared Methods. We compare the detection performance of existing popular methods in the data poisoning literature by adapting them to our setting. Additionally, we include the state-of-the-art influence function method EK-FAC (Grosse et al., 2023) as a baseline. Our $\Delta-\mathrm{Influence}$ method is built upon EK-FAC to ensure fairness in the influence computation itself. This allows us to provide a fair comparison of our algorithm against naive influence function.
@@
    EK-FAC (Grosse et al., 2023) serves as our baseline method for using influence functions in poison detection. It calculates influence scores for every training sample based on one known affected test sample. Samples with average influence scores exceeding a predefined threshold are removed.
Metrics. We evaluate our algorithm using four key metrics. All metrics are averaged over three runs with different random seeds.
@@
    True Positive Rate: Fraction of identified poisoned samples out of the total poisoned samples in train set.
    Number of correctly flagged poisoned samplesTotal number of poisoned $\footnotesize\frac{\text{Number of correctly flagged poisoned samples}}{\text{Total number of poisoned samples}}\times 100\%$
@@
    Precision: Proportion of correctly identified poisoned samples among all samples flagged as poisoned. It reflects the trade-off between detection accuracy and model utility.
    Number of correctly flagged poisoned samplesTotal number of samples flagged as $\footnotesize\frac{\text{Number of correctly flagged poisoned samples}}{\text{Total number of samples flagged as poisoned}}\times 100\%$
@@
    Poison Success Rate: Fraction of poisoned test samples that are misclassified into the target (incorrect) class. For the Witches’ Brew attack, it indicates whether the test point remains poisoned.
    Number of poisoned samples classified as targetTotal number of poisoned $\footnotesize\frac{\text{Number of poisoned samples classified as target}}{\text{Total number of poisoned samples}}\times 100\%$
@@
    Test Accuracy: The Model’s performance on unpoisoned test samples, measuring drop in model utility.
    Number of correct predictions on test setTotal number of test $\footnotesize\frac{\text{Number of correct predictions on test set}}{\text{Total number of test samples}}\times 100\%$
@@
BadNet Frequency Witches’ BadNet Frequency Witches’
@@
SpecSig Precision 3.6% 1.3% 1.4% 1.3% 0.5% 0.3%
TPR 88.3% 88.3% 96.8% 82.6% 78.4% 35.2%
ActClust Precision 2.2% 2.2% 2.1% 1.6% 0.6% 0.3%
TPR 94.9% 99.1% 93.4% 96.3% 100% 55.2%
FreqDef Precision 8.0% 0.4% 10.2% 5.3% 0.1% 1.8%
TPR 72.3% 3.2% 93.6% 85.7% 2.4% 78.4%
EK-FAC Precision 2.8% 2.9% 0.8% 3.2% 0.9% 0.4%
TPR 67.1% 100% 17.4% 70.0% 96.8% 47.2%
$\Delta$-Infl (Ours) Precision 17.6% 13.3% 3.3% 37.3% 2.9% 2.1%
TPR 99.1% 100% 19.4% 96.9% 100% 62.4%
Table 1: Comparison of Precision & TPR across methods and dataset for detecting poisoned samples. Green indicates successful unlearning (poisoning success rate $\leq 5\%$, while red indicates unsuccessful unlearning (see Figure 3 for exact poisoning success rates). We evaluate the precision and true positive rate (TPR) of detecting poisoned training samples. SpecSig (Tran et al., 2018), ActClust (Chen et al., 2018), and EK-FAC (Grosse et al., 2023) yield low precision, flagging many clean samples as poisoned. FreqDef (Zeng et al., 2021) and $\Delta-\mathrm{Influence}$ (Ours) better preserve clean data, though FreqDef shows a significantly lower TPR, missing many true poisoned samples.

### v2

@@ Abstract
Addressing data integrity challenges, such as unlearning the effects of data poisoning after model training, is necessary for the reliable deployment of machine learning models. State-of-the-art influence functions, such as EK-FAC (Grosse et al., 2023) and TRAK (Park et al., 2023), often fail to accurately attribute abnormal model behavior to specific poisoned training data responsible for the data poisoning attack. In addition, traditional unlearning algorithms often struggle to effectively remove the influence of poisoned samples (Pawelczyk et al., 2024), particularly when only a few affected examples can be identified (Goel et al., 2024). To address these challenges, we introduce $\Delta$-Influence, a novel approach that leverages influence functions to trace abnormal model behavior back to the responsible poisoned training data using just one poisoned test example, without assuming any prior knowledge of the attack. $\Delta$-Influence applies data transformations that sever the link between poisoned training data and compromised test points without significantly affecting clean data. This allows detecting large negative shifts in influence scores following data transformations, a phenomenon we term as influence collapse, thereby accurately identifying poisoned training data. Unlearning this subset, e.g. through retraining, effectively eliminates the data poisoning. We validate our method across three vision-based poisoning attacks and three datasets, benchmarking against five detection algorithms and five unlearning strategies. We show that $\Delta$-Influence consistently achieves the best unlearning across all settings, showing the promise of influence functions for corrective unlearning. Our code is publicly available at: https://github.com/Ruby-a07/delta-influence
@@ 1 Introduction
Machine learning models are increasingly deployed in critical sectors such as healthcare, finance, and autonomous systems (Davenport and Kalakota, 2019; Huang et al., 2020; Chatila et al., 2021; Soori et al., 2023). This underscores the importance of ensuring model integrity and robustness against data poisoning attacks. In data poisoning, adversaries intentionally manipulate training data by introducing carefully crafted, often imperceptible modifications (Chatila et al., 2021), leading to incorrect predictions or embedding specific malicious behaviors within the trained models (Fan et al., 2022). Given the large scale of modern datasets, identifying and removing all manipulated samples is typically impractical (Nguyen et al., 2024a; Goel et al., 2024). Therefore, a viable approach involves detecting and attributing the impact of data poisoning to a small set of influential training data points, which is unlearned to mitigate the data poisoning attack.
@@
The challenge of effective unlearning largely depends on the extent of knowledge about the data poisoning attack. For example, Goel et al. (2024) demonstrate that retraining a model after removing a randomly sampled subset containing half of the manipulated data fails to eliminate poisoning in relatively simple attacks like BadNet (Gu et al., 2019). In contrast, retraining without the entire set of manipulated data successfully removes the attack. Furthermore, for more sophisticated poisoning strategies such as Witches’ Brew (Geiping et al., 2021), Pawelczyk et al. (2024) reveal that existing unlearning algorithms are ineffective unless the model is retrained without the full manipulated set, even when full access to the manipulated data is available.
Building upon the framework of Corrective Unlearning introduced by Goel et al. (2024), our work addresses the setting in which the defender has identified a small set of affected test points. We note that detecting such affected data is a practical trigger for realizing that unlearning is necessary and thereby initiating the unlearning process, which can be regarded as a form of poisoning forensics: starting from a compromised output, we trace back to the culpable training examples whose removal neutralizes the attack. In practice, such “perpetrators” typically surface through (i) deployment observation of anomalous behavior (e.g., a permission system granting administrative access to an unknown user, a stop sign being misclassified as a minimum speed-limit sign) or (ii) deliberate in-house stress testing (e.g., red‑teaming, white‑hat). A key advantage of $\Delta-\mathrm{Influence}$ is that it requires only the logically unavoidable minimum that at least one affected test point can be identified; other methods, while effective in their respective settings, typically assume a larger identified set (Min et al., 2025; Coalson et al., 2025). Leveraging this poisoned test point, our approach comprises two primary tasks: first, identifying a critical set of manipulated training points responsible for the compromised prediction; and second, applying unlearning algorithms to remove the influence of these points from the model. Our approach departs from prior unlearning works that often presuppose the availability of a “forget set”, a subset of known poisoned training points (Goel et al., 2023; Kurmanji et al., 2023; Foster et al., 2024). We argue this assumption can be challenging to satisfy in practice, as identifying even one culpable training sample can be difficult, if not impossible, especially in complex clean-label attacks like Witches’ Brew (as illustrated in the ‘Attack’ panel of Figure˜1).
@@
Within this framework, influence functions (Koh and Liang, 2017b) serve as a natural tool for attributing model predictions to specific training data points. However, recent studies (Grosse et al., 2023; Nguyen et al., 2024b; Bae et al., 2024) have indicated that state-of-the-art influence functions struggle to accurately identify the manipulated data when used in a naive manner. Our experiments in Section˜3 also corroborate this finding. To address this limitation, we introduce $\Delta-\mathrm{Influence}$, a novel approach that enhances influence functions to reliably identify a critical set of training data points necessary for unlearning data poisoning without compromising model accuracy. Instead of directly calculating each training point’s influence on a poisoned test point, $\Delta-\mathrm{Influence}$ assesses the change in influence scores before and after perturbing the test point through (i) label flipping and (ii) image transformation. As ablation studies in Section˜4 show, label flipping is essential for breaking the association between poisoned data and the affected test point, while image transformations introduces randomness that reduces false positive rates by preserving the influence of benign data.
@@
To assess the effectiveness of $\Delta-\mathrm{Influence}$ and the broader applicability of influence functions in this context, we apply our method to three prominent data poisoning attacks: Frequency Trigger (Zeng et al., 2021), Witches’ Brew (Geiping et al., 2021), and BadNet (Gu et al., 2019). We compare our approach against multiple defenses (Chen et al., 2018; Tran et al., 2018; Zeng et al., 2021; Grosse et al., 2023; Park et al., 2023) that operate with similar or less information about the poisoning than $\Delta-\mathrm{Influence}$. Each attack presents unique challenges for detection and mitigation, as evidenced by the varying performance of existing detection methods across different attacks. Additionally, we conduct experiments using known unlearning algorithms to unlearn the poisoning attack using the identified set (Golatkar et al., 2020; Chundawat et al., 2023). These experiments provide a comprehensive comparison of these unlearning algorithms. For example, gradient ascent-based methods like SCRUB (Kurmanji et al., 2023) and weight deletion methods like SSD (Foster et al., 2024) can effectively unlearn poisoning when the detected set of training poisons is reasonably accurate. However, their resultant accuracy drops significantly if the detected set includes many falsely flagged clean examples. In contrast, methods like EU and CF (Goel et al., 2024) are surprisingly robust to false positives, delivering the best unlearning and accuracy. Overall, our experiments demonstrate that $\Delta-\mathrm{Influence}$ consistently outperforms existing algorithms across all settings, offering a robust defense against sophisticated data poisoning attacks while preserving accuracy.
@@ 2 Using Influence functions to detect poisons
However, our experiments in Section˜3, along with several recent studies (Nguyen et al., 2024b; Bae et al., 2024; Li et al., 2024b), demonstrate that naively applying state-of-the-art influence functions (Grosse et al., 2023) fails to accurately identify poisoned points in deep neural networks. This limitation necessitates the development of a more robust method to effectively utilize influence functions for detecting and unlearning data poisoning.
@@ 2.1 Our Algorithm: $\Delta$-Influence
To address the shortcomings of the naive approach, we introduce $\Delta-\mathrm{Influence}$. The core idea is to monitor the changes in influence scores of training data points when the affected test point undergoes various transformations.
@@
Formally, let $g_{j}$ be a transformation applied to the test point $z_{\text{te}}=\left(x_{\text{te}},y_{\text{te}}\right)$, consisting of pairing the test image with a random label $y_{\text{te}}^{\prime}$ and applying standard data augmentations such as blurring, color jitter and rotating to $x_{\text{te}}$ (see Appendix B.2 for the list of all transformations). Note that we utilize common data augmentation techniques without designing any poison-specific transformations, suggesting the broad applicability of $\Delta$-Influence. We consider such simplicity to be a key strength of our contribution. Then, for each transformation $g_{j}$, we compute the change in influence score as:
@@
This is illustrated in Figure˜2, where $\Delta\mathrm{Infl}({i},{j})$ is consistently negative for poisoned samples across all transformations, whereas it often remains near zero (compared to that of poisons) or shows no clear trend for clean examples. However, Figure˜2 shows that this is not consistently the case for all clean examples (with some values being considerably small), which brings us to the next component.
Boosting Using Multiple Transformations. The above discussion shows that the change in influence score $\Delta\mathrm{Infl}({i},{j})$ can be used as a score function for detecting whether $z_{\text{tr}}^{i}$ is a manipulated training sample. However, this score function is a relatively weak classifier, especially for clean points, as seen in Figure˜2. To overcome this problem, we use classical ideas from bagging and apply multiple transformations $g_{1},\ldots,g_{n_{b}}$ to obtain a series of weak score functions. Specifically, we use $n_{b}$ transformations to obtain $n_{b}$ weak score functions.
@@
Unlearning identified points. Once the set of poisoned training points $\mathcal{P}$ is identified using $\Delta-\mathrm{Influence}$, the next step is to unlearn these points to mitigate the data poisoning attack. We employ several unlearning algorithms (Goel et al., 2023; Chundawat et al., 2023; Kurmanji et al., 2023; Foster et al., 2024) to remove the influence of $\mathcal{P}$ from the trained model $\theta^{\star}$. In practice, the choice of unlearning algorithm may depend on factors such as computational efficiency, scalability, and the specific characteristics of the poisoning attack. In this work, we look at several popular algorithms including retraining from scratch (denoted as EU (Goel et al., 2023)), CF (Goel et al., 2023), SSD (Foster et al., 2024), SCRUB (Kurmanji et al., 2023), and BadT (Chundawat et al., 2023).
@@ 2.2 Full Algorithm
    Influence Score For each training data point $z_{\text{tr}}^{i}\in\mathcal{D}$ and each transformation $g_{j}\in\mathcal{G}$, compute the change in influence score $\Delta\mathrm{Infl}({i},{j})$ as defined in Equation˜2.
@@ 3.1 Experimental Setup
    Frequency Trigger  (Zeng et al., 2021): In this approach, along with changing the label, a trained, imperceptible pattern is embedded in both the spatial and frequency domains, thereby encompassing the whole image. As shown in Alex et al. (2024), these patterns are difficult to detect by both human and automated methods, making the poisoned samples challenging to identify.
@@
    Clean Label Attack (Witches’ Brew)  (Geiping et al., 2021): Unlike Frequency Trigger, this attack adds an imperceptible pattern to images without altering their labels. The poisoned samples appear benign since their labels are consistent with their content, yet they cause the model to learn incorrect associations, leading to misclassifications during inference. As shown in Pawelczyk et al. (2024), these patterns are difficult to unlearn using unlearning algorithms.
@@
    Patch Trigger (BadNet)  (Gu et al., 2019): Also studied in Goel et al. (2024), this attack involves adding a subtle patch to the corner of selected training images and altering their labels to a designated target class. The presence of the patch causes the model to misclassify any test image containing the patch into the target class while maintaining normal performance on other inputs.
Model and Datasets. We utilize the CIFAR10 and CIFAR100 datasets (Krizhevsky, 2009) and a ResNet18 model (He et al., 2015), following the standard benchmarks and models used in the state-of-the-art machine unlearning setup (Pawelczyk et al., 2024). For CIFAR10, we poison 500 training images (1% of the dataset), while for CIFAR100, we poison 125 training images (0.25% of the dataset) for all attack types except BadNet, which requires a higher size of 350 samples to be effective. The victim class and attack class (when different) are selected randomly. Detection methods are tuned on a small validation set using cross-validation techniques. Hyperparameters such as threshold values and clustering parameters are optimized based on validation performance metrics to achieve the best balance between detection accuracy and false positive rates. Detailed hyperparameter settings and our code are provided in the Appendix B to ensure reproducibility.
Compared Methods. We compare the detection performance of existing popular methods in the data poisoning literature by adapting them to our setting. Additionally, we include the state-of-the-art methods for computing influence function: EK-FAC (Grosse et al., 2023) and TRAK (Park et al., 2023) as baselines. Our $\Delta-\mathrm{Influence}$ method is built upon EK-FAC.
@@
    EK-FAC (Grosse et al., 2023) serves as our baseline method for using influence functions in poison detection. It calculates influence scores for every training sample based on one known affected test sample. Samples with average scores exceeding a predefined threshold are removed.
  5. 5.
    TRAK (Park et al., 2023) uses another implementation of influence functions when thresholding.
Metrics. We evaluate our algorithm using five key metrics. All metrics are averaged over three random seeds.
@@
    True Positive Rate (TPR): Fraction of identified poisoned samples out of the total poisoned samples in train set.
[…]
@@
    Precision: Proportion of correctly identified poisoned samples among all flagged samples. It captures the trade-off between detection accuracy and model utility.
[…]
@@
    Poison Success Rate (PSR): Fraction of poisoned test samples that are misclassified into the target (incorrect) class.
[…]
@@
    Test Accuracy: The performance on unpoisoned test samples, measuring drop in model utility.
[…]
  5. 5.
    Area under the ROC curve (AuROC): Trade-off between TPR and Precision due to the choice of threshold.
@@
Frequency Witches’ BadNet Frequency Witches’ BadNet
@@
SpecSig Precision 1.3% 1.4% 3.6% 0.5% 0.3% 1.3%
TPR 88.3% 96.8% 88.3% 78.4% 35.2% 82.6%
AuROC 0.53 0.88 0.82 0.72 0.58 0.76
ActClust Precision 2.2% 2.1% 2.2% 0.6% 0.3% 1.6%
TPR 99.1% 93.4% 94.9% 100% 55.2% 96.3%
AuROC 0.77 0.75 0.79 0.79 0.53 0.78
FreqDef Precision 0.4% 10.2% 8.0% 0.1% 1.8% 5.3%
TPR 3.2% 93.6% 72.3% 2.4% 78.4% 85.7%
AuROC 0.32 0.98 0.97 0.28 0.91 0.97
TRAK Precision 1.4% 1.0% 1.9% 0.5% 0.2% 1.4%
TPR 70.6% 49.8% 93.6% 96.8% 48.0% 100%
AuROC 0.73 0.50 0.60 0.79 0.49 0.58
EK-FAC Precision 2.9% 0.8% 2.8% 0.9% 0.4% 3.2%
TPR 100% 17.4% 67.1% 96.8% 47.2% 70.0%
AuROC 0.89 0.57 0.87 0.94 0.68 0.71
Precision 13.3% 3.3% 17.6% 2.9% 2.1% 37.3%
$\Delta$-Infl (Ours) TPR 100% 19.4% 99.1% 100% 62.4% 96.9%
AuROC 0.96 0.38 0.95 0.96 0.75 0.82
Table 1: Comparison of Precision & TPR & AuROC across methods and dataset for detecting poisoned samples. Green indicates successful unlearning (PSR $\leq 5\%$, while red indicates failed unlearning (see Figure˜3 for exact poisoning success rates). We evaluate the precision and TPR of detecting poisoned training samples. SpecSig Tran et al. (2018), ActClust Chen et al. (2018), TRAK Park et al. (2023) and EK-FAC Grosse et al. (2023) yield low precision, flagging many clean samples as poisoned. FreqDef Zeng et al. (2021) and $\Delta-\mathrm{Influence}$ better preserve clean data, though FreqDef shows a significantly lower TPR, missing many true poisoned samples. For BadNet, the poisoning success rate correlates with the number of detected poisoned samples, making the attack in Goel et al. (2024) relatively easy to unlearn. In contrast, the Frequency attack requires nearly all poisoned samples to be removed for recovery, making it particularly challenging. Surprisingly, the Witches’ Brew setting is easier than anticipated Pawelczyk et al. (2024), requiring only a few key samples—mainly identified by influence functions—to be removed for effective unlearning.